Alpha Graphics Press (AGP) uses the residual dividend approach to determine the company’s dividend payments. They have a target debt-to-equity ratio of ½. Next year AGP plans a capital budget of $60 million and earnings of $50 million. How much will AGP pay in dividends next year? Show your work.
Answer #1

Capital budget=60 million
Equity=2/3*60=40 million
Debt=1/3*60=20 million
Earnings=50 million

So, out of 50 million earnings, 40 million would be used as equity and the remaining 10 million would be paid out as dividends.

Hence, Dividends next year=10 million